OpenStudy (anonymous):

can anyone help me simplify 2n!/(2n-3)! Really haven't gotten the factorial concept yet :(

OpenStudy (anonymous):

2n! / (2n-3)(2n-2)(2n-1)(2n!) Then cancel 2n! 1/(2n-3)(2n-2)(2n-1)

OpenStudy (cwrw238):

2n *( 2n-1) * (2n-2)

OpenStudy (anonymous):

yeah how did u get 2 2n*2n-1*2n-2???? Please explain :)

OpenStudy (anonymous):

Sub x = 2n to make it simpler

OpenStudy (anonymous):

x! / (x-3)! = x (x-1) (x-2)

OpenStudy (anonymous):

This the basic definition of factorial: \[\large n! = n \times (n-1) \times (n-2) ..................1\] Here keep on decreasing 1..

OpenStudy (anonymous):

Ty telliott99 & cwrw238 =) =)

OpenStudy (cwrw238):

compare 6! = 6*5*4*3*2*1 6 - 3 = 3 3! = 3*2*1 6! 6*5*4*3*2*1 -- = ----------- = 6*5*4 3! 3*2*1

OpenStudy (anonymous):

Hope you've got it now.

OpenStudy (anonymous):

yeah got it :D finally!!!